# TensorHyper-VQC: A Tensor-Train-Guided Hypernetwork for Robust and Scalable Variational Quantum Computing

### v3

@@ Abstract
Variational Quantum Computing (VQC) faces fundamental scalability barriers, primarily due to barren plateaus and sensitivity to quantum noise. To address these challenges, we introduce TensorHyper-VQC, a novel tensor-train (TT)-guided hypernetwork framework that significantly improves the robustness and scalability of VQC. Our framework fully delegates the generation of quantum circuit parameters to a classical TT network, effectively decoupling optimization from quantum hardware. This innovative parameterization mitigates gradient vanishing, enhances noise resilience through structured low-rank representations, and facilitates efficient gradient propagation. Grounded in Neural Tangent Kernel and statistical learning theory, our rigorous theoretical analyses establish strong guarantees on approximation capability, optimization stability, and generalization performance. Extensive empirical results across quantum dot classification, Max-Cut optimization, and molecular quantum simulation tasks demonstrate that TensorHyper-VQC consistently achieves superior performance and robust noise tolerance, including hardware-level validation on a 156-qubit IBM Heron processor. These results position TensorHyper-VQC as a scalable and noise-resilient framework for advancing practical quantum machine learning on near-term devices.
@@ 2.1 The TensorHyper-VQC Framework
where each $\mathcal{G}_{k}$ is a low-rank TT-core, and $K$ represents the number of TT-cores in the TT decomposition. This design ensures that gradient computations and parameter updates are solely performed on the TT-cores using standard backpropagation in the classical domain.
@@ 2.2 Theoretical Guarantees
Suppose that the quantum measurement noise is represented by a random vector $\boldsymbol{\tau}\in\mathbb{R}^{3U}$ with mean $\textbf{b}=\mathbb{E}[\boldsymbol{\tau}]$ and covariance matrix $\boldsymbol{\Sigma}=\text{Cov}(\boldsymbol{\tau})$, and given $K_{\ell}$-Lipschitz loss function and dataset $\{(\textbf{x}_{n},\textbf{y}_{n})\}_{n=1}^{N}$, the generalization gap obeys:
$\mathbb{E}\left[\sup\limits_{\boldsymbol{\theta}}\left|\mathcal{R}(f_{\boldsymbol{\theta}})-\hat{\mathcal{R}}(f_{\boldsymbol{\theta}})\right|\right]\leq\tilde{O}\left(\sqrt{\frac{\text{Tr}(\mathcal{T}_{\rm tv})}{N^{2}}}\right)+\mathcal{O}\left(\sqrt{\frac{\lambda_{\rm max}(\boldsymbol{\Sigma})K_{\ell}^{2}}{N}}\right)+\mathcal{O}\left(\frac{\lVert\textbf{b}\rVert_{2}K_{\ell}}{\sqrt{N}}\right),$ (15)
@@
Let $r=\max_{i}r_{i}$. Then increasing $r$ reduces the TT approximation error $\epsilon_{\rm tt}(r)$, and hence improves expressivity, but simultaneously enlarges $\text{Tr}(\mathcal{T}_{\rm tv})$, leading to a larger upper bound on the expected generalization gap $\mathbb{E}\left[\sup_{\boldsymbol{\theta}}\left|\mathcal{R}(f_{\boldsymbol{\theta}})-\hat{\mathcal{R}}(f_{\boldsymbol{\theta}})\right|\right]$.
@@
$\frac{\partial\hat{\mathcal{R}}}{\partial\mathcal{G}_{k}}=\left(\frac{\partial\hat{\mathcal{R}}}{\partial\hat{\textbf{w}}}+\boldsymbol{\tau}\right)\cdot\frac{\partial\hat{\textbf{w}}}{\partial\mathcal{G}_{k}}=\sum\limits_{u=1}^{3UL}\frac{\partial\mathcal{R}}{\partial\hat{w}_{u}}\cdot\frac{\partial\hat{w}_{u}}{\partial\mathcal{G}_{k}}+\sum\limits_{u=1}^{3UL}\tau_{u}\frac{\partial\hat{w}_{u}}{\partial\mathcal{G}_{k}}.$ (18)
@@
Furthermore, if we assume that the quantum measurement noise is $\boldsymbol{\tau}\in\mathbb{R}^{3UL}$ with mean $\textbf{b}:=\mathbb{E}[\boldsymbol{\tau}]$ and covariance $\boldsymbol{\Sigma}=\text{Cov}(\boldsymbol{\tau})$. With $\textbf{J}_{k}:=\frac{\partial\hat{\textbf{w}}}{\partial\mathcal{G}_{k}}$, the TT-core gradient estimator is
@@
$\mathbb{E}[\hat{\textbf{g}}_{k}]-\textbf{g}_{k}=\textbf{J}_{k}^{\top}\textbf{b},\hskip 11.38109pt\text{Var}(\hat{\textbf{g}}_{k})=\textbf{J}_{k}^{\top}\boldsymbol{\Sigma}\textbf{J}_{k},$ (25)
@@
$\left\lVert\mathbb{E}[\hat{\textbf{g}}_{k}]-\textbf{g}_{k}\right\rVert_{2}\leq\lVert\textbf{J}_{k}\rVert_{2}\lVert\textbf{b}\rVert_{2}\leq\sqrt{\frac{c}{3UL}}\lVert\textbf{b}\rVert_{2},$ (26)
@@
$\text{Var}(\hat{\textbf{g}}_{\textbf{w}})\approx\lambda_{\rm max}(\boldsymbol{\Sigma}),\hskip 11.38109pt\mathbb{E}[\hat{\textbf{g}}_{\textbf{w}}]-\textbf{g}_{\textbf{w}}=\textbf{b},$ (30)
@@ 2.4 Application to Quantum Machine Learning: Quantum Dot Classification
Performance in the Quantum Noise-Free Regime. We first benchmark TensorHyper-VQC against standard VQC, MLPHyper-VQC, and the hybrid TTN+VQC model in a noise-free setting, as shown in Fig. 3(a). TensorHyper-VQC rapidly converges to near-perfect accuracy ($99.5\%$) within $10$ epochs, significantly outperforming both standard VQC ($65.7\%$) and TTN+VQC ($91.7\%$), as well as the MLPHyper-VQC baseline (56.5%). Compared with deep classical models (ResNet18+LoRA and ResNet50+LoRA), TensorHyper-VQC also achieves higher accuracy with 46x4-6x fewer parameters, illustrated in Fig. 3(b). These results empirically confirm that the TensorHyper-VQC architecture offers a distinct advantage in terms of expressivity and optimization efficiency compared to conventional quantum, hybrid, and classical approaches.
@@ 2.5 Application to Combinatorial Optimization: Max-Cut
With Pauli-Z operators $\sigma_{z}^{(i)}$ on qubit $i$, the Max-Cut cost function is encoded as:
$\mathcal{H}=\sum\limits_{(i,j)\in E}\frac{1-\sigma_{z}^{(i)}\sigma_{z}^{(j)}}{2}.$ (34)
In our experiments, TensorHyper-VQC employs a compact TT network to generate all variational circuit parameters required by QAOA across different graph instances. The TT network is constructed with an input dimension of $10$, output dimensions $[1,1]$, and TT-ranks $[1,4,1]$, resulting in a lightweight parameterization with only $41$ trainable parameters (including bias). We benchmarked the approach on 10 randomly generated 20-qubit Erdos-Renyi graphs [12], where each edge is included independently with probability $0.5$.
@@ 2.6 Application to Quantum Simulation: Molecular Hamiltonians
To further demonstrate the practical utility of the proposed TensorHyper-VQC framework, we evaluated its performance on a canonical quantum simulation task: the ground-state energy estimation of the lithium hydride (LiH) molecule. Quantum chemistry simulation is widely regarded as one of the most compelling near-term applications of variational quantum algorithms. It thus serves as a stringent benchmark for assessing the scalability, expressivity, and robustness of variational architectures. In this study, we considered a 4-qubit reduced Hamiltonian of LiH, obtained by freezing core orbitals and truncating high-lying virtual orbitals on an STO-3G basis. This reduced model preserves the essential many-body correlations while enabling tractable simulations on current quantum hardware.
@@ 3 Discussion
Conventional VQCs often suffer from vanishing gradients and noise-induced instability, limiting their expressivity and applicability to realistic quantum devices. Numerous recent studies have attempted to address these challenges through advanced parameterization strategies [23] and structural optimizations [16]. Low-rank tensor decompositions, in particular, have emerged as a practical approach to reduce parameter dimensionality and enhance generalization in both classical and quantum settings. Building on these advances, TensorHyper-VQC leverages an entirely classical TT hypernetwork to generate variational parameters, thereby combining the representational power of tensor networks with the trainability benefits of low-rank optimization.
@@ 4.1 An Introduction to Neural Tangent Kernel Theory
The Neural Tangent Kernel (NTK) [3, 19, 25] is a potent theoretical framework that illustrates the efficient trainability of deep neural networks despite their inherent complexity. Fundamentally, the NTK theory examines how marginal changes in a network’s parameters affect its predictions. A network is deemed ”well-conditioned,” leading to smoother training and faster convergence, when minor parameter adjustments reliably yield stable, predictable output shifts.
@@ 4.2 Tensor-Train Network and TTN+VQC
When integrated with VQC, the TT network forms a hybrid quantum-classical architecture, traditionally referred to as TTN+VQC [35]. In this framework, the TT network typically serves as a classical pre-processing or feature-compression module that reduces the practical input dimension before quantum encoding. The VQC then processes the compressed features, leveraging quantum superposition and entanglement to enhance the model’s expressivity and capture nonlinear correlations that are difficult to represent classically. This complementary interplay positions TTN+VQC as a promising paradigm for tackling complex and high-dimensional quantum machine learning tasks.
@@ 4.3 Quantum Approximate Optimization Algorithm
The Quantum Approximate Optimization Algorithm (QAOA) [49, 50] is a hybrid quantum-classical algorithm designed to find approximate solutions to combinatorial optimization problems. It iteratively optimizes the parameters of a quantum circuit to maximize an objective function defined by the problem. One common application of QAOA is solving the Max-Cut problem. The Max-Cut problem seeks to partition the vertices of a graph into two sets, thereby maximizing the number of edges between the two sets.
@@ 4.9 Proof of Upper Bound on Generalization Performance
$\mathbb{E}\left[\sup\limits_{f_{\boldsymbol{\theta}}\in\mathcal{F}_{\boldsymbol{\theta}}}|\mathcal{R}(f_{\boldsymbol{\theta}})-\hat{\mathcal{R}}(f_{\boldsymbol{\theta}})|\right]\leq 2\mathcal{C}_{N}(\mathcal{F}_{\boldsymbol{\theta}})+\left(\text{Var}\left[\frac{\partial\mathcal{R}(f_{\boldsymbol{\theta}})}{\partial\boldsymbol{\theta}}\right]\right)^{\frac{1}{2}}.$ (46)
@@
On the other hand, suppose the quantum measurement noise is represented by a random vector $\boldsymbol{\tau}\in\mathbb{R}^{3U}$ with mean $\textbf{b}=\mathbb{E}[\boldsymbol{\tau}]$ and covariance matrix $\boldsymbol{\Sigma}=\text{Cov}(\boldsymbol{\tau})$. Given a $K_{\ell}$-Lipschitz loss function and a training set $\{(\textbf{x}_{n},\textbf{y}_{n})\}_{n=1}^{N}$, using the linearized NTK approximation, we have:
@@
$\mathbb{E}\left[\sup\limits_{\boldsymbol{\theta}}\left|\mathcal{R}(f_{\boldsymbol{\theta}})-\hat{\mathcal{R}}(f_{\boldsymbol{\theta}})\right|\right]\leq\tilde{O}\left(\sqrt{\frac{\text{Tr}(\mathcal{T}_{\rm tv})}{N^{2}}}\right)+\mathcal{O}\left(\sqrt{\frac{\lambda_{\rm max}(\boldsymbol{\Sigma})K_{\ell}^{2}}{N}}\right)+\mathcal{O}\left(\frac{\lVert\textbf{b}\rVert_{2}K_{\ell}}{\sqrt{N}}\right),$ (49)
@@
$f_{\boldsymbol{\theta}}(\textbf{x})=\text{VQC}\left(\hat{\textbf{w}}(\textbf{x};\boldsymbol{\theta})\right),$ (52)
@@ 4.10 Proof of Robustness to Quantum Noise
We first consider the issue of noise propagation in gradient computation. Let $\hat{\textbf{w}}\in\mathbb{R}^{3UL}$ denote all quantum gate parameters and $\{\mathcal{G}_{k}\}_{k=1}^{K}$ the TT cores. The Jacobian $\textbf{J}_{k}$ is defined as:
@@
$\mathbb{E}[\hat{\textbf{g}}_{k}]-\textbf{g}_{k}=\textbf{J}_{k}^{\top}\textbf{b},\hskip 7.11317pt\text{Var}(\hat{\textbf{g}}_{k})\lesssim\frac{c\lambda_{\rm max}(\boldsymbol{\Sigma})}{3UL}.$ (64)
@@
$\text{Var}(\hat{\textbf{g}}_{\textbf{w}})=\boldsymbol{\Sigma},\hskip 8.53581pt\mathbb{E}[\hat{\textbf{g}}_{\textbf{w}}]-\textbf{g}_{\textbf{w}}=\textbf{b},$ (66)

### v4

@@ Abstract
Variational Quantum Computing (VQC) faces fundamental scalability barriers, primarily due to barren plateaus and sensitivity to quantum noise. To address these challenges, we introduce TensorHyper-VQC, a novel tensor-train (TT)-guided hypernetwork framework that significantly improves the robustness and scalability of VQC. Our framework fully delegates the generation of quantum-circuit parameters to a classical TT network, thereby decoupling optimization from quantum hardware. This innovative parameterization mitigates gradient vanishing, enhances noise resilience through structured low-rank representations, and facilitates efficient gradient propagation. Grounded in Neural Tangent Kernel and statistical learning theory, our rigorous theoretical analyses establish strong guarantees on approximation capability, optimization stability, and generalization performance. Extensive empirical results across quantum dot classification, Max-Cut optimization, and molecular quantum simulation tasks demonstrate that TensorHyper-VQC consistently achieves superior performance and robust noise tolerance, including hardware-level validation on a 156-qubit IBM Heron processor. These results position TensorHyper-VQC as a scalable and noise-resilient framework for advancing practical quantum machine learning on near-term devices.
@@ 2.1 The TensorHyper-VQC Framework
where each $\mathcal{G}_{k}$ is a low-rank TT-core, and $K$ represents the number of TT-cores in the TT decomposition. This design ensures that gradient computations and parameter updates are performed solely on the TT-cores, using standard backpropagation in the classical domain.
@@ 2.2 Theoretical Guarantees
Suppose that the quantum measurement noise is represented by a random vector $\boldsymbol{\tau}\in\mathbb{R}^{3U}$ with mean $\textbf{b}=\operatorname{E}[\boldsymbol{\tau}]$ and covariance matrix $\boldsymbol{\Sigma}=\text{Cov}(\boldsymbol{\tau})$, and given $K_{\ell}$-Lipschitz loss function and dataset $\{(\textbf{x}_{n},\textbf{y}_{n})\}_{n=1}^{N}$, the generalization gap obeys:
$\operatorname{E}\left[\sup\limits_{\boldsymbol{\theta}}\left|\mathcal{R}(f_{\boldsymbol{\theta}})-\hat{\mathcal{R}}(f_{\boldsymbol{\theta}})\right|\right]\leq\tilde{O}\left(\sqrt{\frac{\text{Tr}(\mathcal{T}_{\rm tv})}{N^{2}}}\right)+\mathcal{O}\left(\sqrt{\frac{\lambda_{\rm max}(\boldsymbol{\Sigma})K_{\ell}^{2}}{N}}\right)+\mathcal{O}\left(\frac{\lVert\textbf{b}\rVert_{2}K_{\ell}}{\sqrt{N}}\right),$ (15)
@@
Let $r=\max_{i}r_{i}$. Then increasing $r$ reduces the TT approximation error $\epsilon_{\rm tt}(r)$, and hence improves expressivity, but simultaneously enlarges $\text{Tr}(\mathcal{T}_{\rm tv})$, leading to a larger upper bound on the expected generalization gap $\operatorname{E}\left[\sup_{\boldsymbol{\theta}}\left|\mathcal{R}(f_{\boldsymbol{\theta}})-\hat{\mathcal{R}}(f_{\boldsymbol{\theta}})\right|\right]$.
@@
$\frac{\partial\hat{\mathcal{R}}}{\partial\mathcal{G}_{k}}=\left(\frac{\partial\mathcal{R}}{\partial\hat{\textbf{w}}}+\boldsymbol{\tau}\right)\cdot\frac{\partial\hat{\textbf{w}}}{\partial\mathcal{G}_{k}}=\sum\limits_{u=1}^{3UL}\frac{\partial\mathcal{R}}{\partial\hat{w}_{u}}\cdot\frac{\partial\hat{w}_{u}}{\partial\mathcal{G}_{k}}+\sum\limits_{u=1}^{3UL}\tau_{u}\frac{\partial\hat{w}_{u}}{\partial\mathcal{G}_{k}}.$ (18)
@@
Furthermore, if we assume that the quantum measurement noise is $\boldsymbol{\tau}\in\mathbb{R}^{3UL}$ with mean $\textbf{b}:=\operatorname{E}[\boldsymbol{\tau}]$ and covariance $\boldsymbol{\Sigma}=\text{Cov}(\boldsymbol{\tau})$. With $\textbf{J}_{k}:=\frac{\partial\hat{\textbf{w}}}{\partial\mathcal{G}_{k}}$, the TT-core gradient estimator is
@@
$\operatorname{E}[\hat{\textbf{g}}_{k}]-\textbf{g}_{k}=\textbf{J}_{k}^{\top}\textbf{b},\hskip 11.38109pt\text{Var}(\hat{\textbf{g}}_{k})=\textbf{J}_{k}^{\top}\boldsymbol{\Sigma}\textbf{J}_{k},$ (25)
@@
$\left\lVert\operatorname{E}[\hat{\textbf{g}}_{k}]-\textbf{g}_{k}\right\rVert_{2}\leq\lVert\textbf{J}_{k}\rVert_{2}\lVert\textbf{b}\rVert_{2}\leq\sqrt{\frac{c}{3UL}}\lVert\textbf{b}\rVert_{2},$ (26)
@@
$\text{Var}(\hat{\textbf{g}}_{\textbf{w}})\approx\lambda_{\rm max}(\boldsymbol{\Sigma}),\hskip 11.38109pt\operatorname{E}[\hat{\textbf{g}}_{\textbf{w}}]-\textbf{g}_{\textbf{w}}=\textbf{b},$ (30)
@@ 2.4 Application to Quantum Machine Learning: Quantum Dot Classification
Performance in the Quantum Noise-Free Regime. We first benchmark TensorHyper-VQC against standard VQC, MLPHyper-VQC, and the hybrid TTN+VQC model in a noise-free setting, as shown in Fig. 3(a). TensorHyper-VQC rapidly converges to near-perfect accuracy ($99.5\%$) within $10$ epochs, significantly outperforming both standard VQC ($65.7\%$) and TTN+VQC ($91.7\%$), as well as the MLPHyper-VQC baseline (56.5%). Compared with deep classical models (ResNet18+LoRA and ResNet50+LoRA), TensorHyper-VQC also achieves higher accuracy with 46x4-6x fewer parameters, illustrated in Fig. 3(b). These results empirically confirm that the TensorHyper-VQC architecture offers a distinct advantage in expressivity and optimization efficiency compared with conventional quantum, hybrid, and classical approaches.
@@ 2.5 Application to Combinatorial Optimization: Max-Cut
With Pauli-Z operators $\sigma_{z}^{(i)}$, $\sigma_{z}^{(j)}$ on qubit $i$ and $j$ associated with the cost $c_{ij}$, the Max-Cut cost function is encoded as:
$\mathcal{H}=\sum\limits_{(i,j)\in E}c_{ij}\frac{1-\sigma_{z}^{(i)}\sigma_{z}^{(j)}}{2}.$ (34)
In our experiments, TensorHyper-VQC employs a compact TT network to generate all variational circuit parameters required by QAOA across different graph instances. The TT network is constructed with an input dimension of $10$, output dimensions $[1,1]$, and TT-ranks $[1,4,1]$, resulting in a lightweight parameterization with only $41$ trainable parameters (including bias). We benchmarked the approach on 10 randomly generated 20-qubit Erdos-Renyi graphs [12], in which each edge is included independently with probability $0.5$.
@@ 2.6 Application to Quantum Simulation: Molecular Hamiltonians
To further demonstrate the practical utility of the proposed TensorHyper-VQC framework, we evaluated its performance on a canonical quantum simulation task: the ground-state energy estimation of the lithium hydride (LiH) molecule. Quantum chemistry simulations are widely regarded as among the most compelling near-term applications of variational quantum algorithms. It thus serves as a stringent benchmark for assessing the scalability, expressivity, and robustness of variational architectures. In this study, we considered a 4-qubit reduced Hamiltonian of LiH, obtained by freezing core orbitals and truncating high-lying virtual orbitals on an STO-3G basis. This reduced model preserves the essential many-body correlations while enabling tractable simulations on current quantum hardware.
@@ 3 Discussion
Conventional VQCs often suffer from vanishing gradients and noise-induced instability, limiting their expressivity and applicability to realistic quantum devices. Numerous recent studies have sought to address these challenges through advanced parameterization strategies [23] and structural optimization [16]. Low-rank tensor decompositions, in particular, have emerged as a practical approach to reduce parameter dimensionality and enhance generalization in both classical and quantum settings. Building on these advances, TensorHyper-VQC leverages an entirely classical TT hypernetwork to generate variational parameters, thereby combining the representational power of tensor networks with the trainability benefits of low-rank optimization.
@@ 4.1 An Introduction to Neural Tangent Kernel Theory
The Neural Tangent Kernel (NTK) [3, 19, 25] is a powerful theoretical framework that explains the efficient learnability of deep neural networks despite their inherent complexity. Fundamentally, the NTK theory examines how marginal changes in a network’s parameters affect its predictions. A network is deemed ”well-conditioned,” leading to smoother training and faster convergence, when minor parameter adjustments reliably yield stable, predictable output shifts.
@@ 4.2 Tensor-Train Network and TTN+VQC
When integrated with VQC, the TT network forms a hybrid quantum-classical architecture, traditionally referred to as TTN+VQC [35]. In this framework, the TT network typically serves as a classical preprocessing or feature-compression module that reduces the input dimensionality before quantum encoding. The VQC then processes the compressed features, leveraging quantum superposition and entanglement to enhance the model’s expressivity and capture nonlinear correlations that are difficult to represent classically. This complementary interplay positions TTN+VQC as a promising paradigm for tackling complex and high-dimensional quantum machine learning tasks.
@@ 4.3 Quantum Approximate Optimization Algorithm
The Quantum Approximate Optimization Algorithm (QAOA) [49, 50] is a hybrid quantum-classical algorithm designed to find approximate solutions to combinatorial optimization problems. It iteratively optimizes the parameters of a quantum circuit to maximize an objective function defined by the problem. One common application of QAOA is solving the Max-Cut problem. The Max-Cut problem seeks to partition the vertices of a graph into two sets, thereby maximizing the number of edges between them.
@@ 4.9 Proof of Upper Bound on Generalization Performance
$\operatorname{E}\left[\sup\limits_{f_{\boldsymbol{\theta}}\in\mathcal{F}_{\boldsymbol{\theta}}}|\mathcal{R}(f_{\boldsymbol{\theta}})-\hat{\mathcal{R}}(f_{\boldsymbol{\theta}})|\right]\leq 2\mathcal{C}_{N}(\mathcal{F}_{\boldsymbol{\theta}})+\left(\text{Var}\left[\frac{\partial\mathcal{R}(f_{\boldsymbol{\theta}})}{\partial\boldsymbol{\theta}}\right]\right)^{\frac{1}{2}}.$ (46)
@@
On the other hand, suppose the quantum measurement noise is represented by a random vector $\boldsymbol{\tau}\in\mathbb{R}^{3U}$ with mean $\textbf{b}=\operatorname{E}[\boldsymbol{\tau}]$ and covariance matrix $\boldsymbol{\Sigma}=\text{Cov}(\boldsymbol{\tau})$. Given a $K_{\ell}$-Lipschitz loss function and a training set $\{(\textbf{x}_{n},\textbf{y}_{n})\}_{n=1}^{N}$, using the linearized NTK approximation, we have:
@@
$\operatorname{E}\left[\sup\limits_{\boldsymbol{\theta}}\left|\mathcal{R}(f_{\boldsymbol{\theta}})-\hat{\mathcal{R}}(f_{\boldsymbol{\theta}})\right|\right]\leq\tilde{O}\left(\sqrt{\frac{\text{Tr}(\mathcal{T}_{\rm tv})}{N^{2}}}\right)+\mathcal{O}\left(\sqrt{\frac{\lambda_{\rm max}(\boldsymbol{\Sigma})K_{\ell}^{2}}{N}}\right)+\mathcal{O}\left(\frac{\lVert\textbf{b}\rVert_{2}K_{\ell}}{\sqrt{N}}\right),$ (49)
@@
$f_{\boldsymbol{\theta}}(\textbf{x})=\operatorname{VQC}\left(\hat{\textbf{w}}(\textbf{x};\boldsymbol{\theta})\right),$ (52)
@@ 4.10 Proof of Robustness to Quantum Noise
We first consider noise propagation in gradient computation. Let $\hat{\textbf{w}}\in\mathbb{R}^{3UL}$ denote all quantum gate parameters and $\{\mathcal{G}_{k}\}_{k=1}^{K}$ the TT cores. The Jacobian $\textbf{J}_{k}$ is defined as:
@@
$\operatorname{E}[\hat{\textbf{g}}_{k}]-\textbf{g}_{k}=\textbf{J}_{k}^{\top}\textbf{b},\hskip 7.11317pt\text{Var}(\hat{\textbf{g}}_{k})\lesssim\frac{c\lambda_{\rm max}(\boldsymbol{\Sigma})}{3UL}.$ (64)
@@
$\text{Var}(\hat{\textbf{g}}_{\textbf{w}})=\boldsymbol{\Sigma},\hskip 8.53581pt\operatorname{E}[\hat{\textbf{g}}_{\textbf{w}}]-\textbf{g}_{\textbf{w}}=\textbf{b},$ (66)